# VFScale: Intrinsic Reasoning through Verifier-Free Test-time Scalable Diffusion Model

### v3

@@ Abstract
Inspired by human SYSTEM 2 thinking, LLMs excel at complex reasoning tasks via extended Chain-of-Thought. However, similar test-time scaling for diffusion models to tackle complex reasoning remains largely unexplored. From existing work, two primary challenges emerge in this setting: (i) the dependence on an external verifier indicating a notable gap from intrinsic reasoning of human intelligence without any external feedback, and (ii) the lack of an efficient search algorithm. In this paper, we introduce the Verifier-free Test-time Scalable Diffusion Model (VFScale) to achieve scalable intrinsic reasoning, which equips number-of-sample test-time scaling with the intrinsic energy function of diffusion models as the verifier. Concretely, VFScale comprises two key innovations to address the aforementioned challenges. On the training side, VFScale consists of a novel LRNCL loss and a KL regularization to improve the energy landscape, ensuring that the learned energy function itself serves as a reliable verifier. On the inference side, VFScale integrates the denoising process with a novel hybrid Monte Carlo Tree Search (hMCTS) to improve search efficiency. On challenging reasoning tasks of Maze and Sudoku, we demonstrate the effectiveness of VFScale’s training objective and scalable inference method. In particular, trained with Maze sizes of up to $6\times 6$, our VFScale solves 88% of Maze problems with much larger sizes of $15\times 15$, while standard diffusion model completely fails.
@@ 1 Introduction
Human intelligence can allocate more computation to solve harder problems, known as SYSTEM 2 thinking kahneman2011thinking . Motivated by this, large language models (LLMs) have exhibited promising performance in reasoning tasks through deliberate long Chain-of-Thought lightman2023let ; deepseekai2025deepseekr1incentivizingreasoningcapability . Alongside LLMs, diffusion models have recently emerged as a compelling alternative. By casting reasoning as an optimization problem, these models iteratively refine candidate solutions toward higher-quality outputs. This approach has demonstrated strong potential, achieving notable results on tasks such as Sudoku solving, graph connectivity, and shortest-path computation du2022learning ; du2024learning .
@@
Although the iterative denoising procedure of diffusion models mirrors the repeated reasoning cycles required by complex reasoning tasks, their performance declines sharply once the problem difficulty exceeds the training distribution du2022learning . Moreover, prior works show that simply increasing the number of sampling steps quickly leads to a performance plateau du2024learning . Instead, in image generation, inference-time scaling to improve performance by increasing the number of samples has proven effective ma2025inferencetimescalingdiffusionmodels . However, this strategy depends on an external verifier that is often unavailable or prohibitively expensive ouyang2022training ; lee2024rlaif . Yet human intelligence can engage in deep, reflective intrinsic reasoning without any external feedback, revealing a clear gap with such methods lombrozo2024learning ; chi1994eliciting ; huanglarge . These observations motivate our central question: Can we design a Verifier-free Test-time Scalable Diffusion Model to achieve scalable intrinsic reasoning?
@@
In pursuit of such a test-time scalable intrinsic reasoning model, we leverage the fundamental mechanics of diffusion models. Diffusion models approximate the score function song2020score , which is equivalent to the negative gradient of the energy function. Our key insight is that the energy function, as a measure of the learned probability, can serve as an effective verifier itself. Based on the above premise, in this work, we explore equipping number-of-sample test-time scaling with diffusion models’ intrinsic energy function.
@@
However, this strategy presents two primary challenges: (1) Firstly, using number-of-sample test-time scaling requires a verifier to evaluate and select samples. Prior work on test-time scaling has also established that verifier accuracy is crucial for performance setlur2025scalingtesttimecomputeverification ; ma2025inferencetimescalingdiffusionmodels . Consequently, using the energy function as an intrinsic verifier requires it to reliably reflect the sample quality. Nevertheless, previous energy-based diffusion models have largely ignored this requirement. (2) Secondly, the efficiency of the search algorithm is critical when scaling by varying the number of samples snell2024scalingllmtesttimecompute . Nevertheless, the search efficiency of existing methods, such as best-of-$N$ (BoN), warrants further enhancement.
@@
To approach these challenges, we propose Verifier-Free Test-Time Scalable Diffusion Models (VFScale), a novel framework designed to solve complex reasoning problems. VFScale consists of innovations in both training objectives and inference scaling as shown in Fig. 2 to respectively address the two aforementioned challenges. (1) On the training side, VFScale incorporates a novel Linear-Regression Negative Contrastive Learning (LRNCL) objective and a KL regularization to improve the energy landscape, especially to better align the energy function with sample quality with LRNCL loss. The LRNCL objective samples two negatives [^1] Here, positive samples refer to data points, while negative samples correspond to noise-perturbed versions of the data. for each positive sample and fits a linear regression between their energy values and their L2 distances to the positive. It encourages a positively sloped linear relationship, aligning the energy landscape with distance to the ground truth. We refer to this alignment as performance-energy consistency. [^2] Detailed description for performance-energy consistency can be found in Appendix A. The KL regularization term backpropagates through the denoising process, smoothing the energy landscape. (2) On the inference side, we integrate the denoising process with a novel hybrid Monte Carlo Tree Search (hMCTS). It sequentially performs BoN and MCTS as the denoising proceeds. In the initial steps, BoN carries out parallel denoising without selecting to avoid premature exclusion of potential solutions, while the later MCTS denoising allows better navigation in the search space. Taken together, the hybrid MCTS denoising allows the model to significantly improve inference performance by scaling up the inference computational budget.
@@
In summary, our contributions are as follows: (1) We introduce Verifier-free Test-time Scalable Diffusion Models (VFScale), a novel framework that can scale up test-time computation for better reasoning capability. (2) We introduce Linear-Regression Negative Contrastive Learning (LRNCL) and KL regularization into the training objective to improve the energy landscape. (3) We integrate a hybrid Monte Carlo Tree Search into the denoising process, enabling test-time scaling with better search efficiency. (4) We demonstrate the effectiveness of our method on challenging Sudoku and Maze datasets where the conditions are out-of-distribution or the problem sizes are much larger.
Figure 2: Overview of VFScale. This figure illustrates the key aspects of VFScale by contrasting its training and inference strategies with those of the previous method. (1) To qualify the intrinsic energy of diffusion models as a verifier, VFScale introduces ${\mathcal{L}}_{\text{LRNCL}}$ and ${\mathcal{L}}_{\text{KL}}$ to improve the energy landscape during training. (2) In order for a higher search efficiency, VFScale proposes hybrid Monte Carlo Tree Search (hMCTS) that achieves a balance between best-of-$N$ and MCTS.
@@ 2 Related Work
Diffusion for Reasoning: The iterative refinement process inherent in diffusion models lends itself well to tasks requiring multi-step reasoning. du2022learning firstly formulated multi-step reasoning as an energy minimization problem. To approach harder problems with more complex energy landscapes, more compute budget is used for optimization steps. du2024learning formulated the energy-based model as a diffusion model, and the performance on harder tasks was significantly improved. While effective, the performance gain by scaling the compute of these methods quickly saturates as the number of sampling steps increases. Building upon du2024learning , VFScale instead explores number-of-sample scaling and proposes corresponding training and inference strategies.
Test-time Scaling of Diffusion: Improving diffusion model samples by increasing compute resources at inference time has emerged as an increasingly important direction. Early approaches focused on increasing the number of sampling steps, which leads to diminishing returns ho2020denoising ; karras2022elucidating . More recent work ma2025inferencetimescalingdiffusionmodels ; yoon2025montecarlotreediffusion has explored alternative scaling dimensions, such as increasing the number of sampled candidates, though these methods typically rely on external verifiers to guide solution selection. VFScale advances beyond these methods by leveraging the intrinsic energy function to guide test-time search, a capability further cultivated through customized training and inference strategies. These innovations make VFScale a versatile and scalable solution for real-world reasoning tasks.
@@ 3 Preliminary
The Denoising Diffusion Probabilistic Model (DDPM) ho2020denoising contains a predefined forward process to corrupt data into Gaussian noises, and a learnable reverse process to generate new samples from them. The forward process follows a Gaussian transition kernel $q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t}|\sqrt{\alpha_{t}}\bm{x}_{t-1},(1-\alpha_{t})\bm{I}),t=1,\cdots,T$, where the noise schedule $\{\alpha_{t}\}_{t=1}^{T}$ and $T$ is set to make $\bm{x}_{T}$ follow approximately a standard Gaussian distribution. The reverse process can be learned to predict the noise from the corrupted data by minimizing
@@
where the expectation is w.r.t. $\bm{x}_{0}\sim p(\bm{x})$, $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$, and $t\sim\{1,...T\}$. $\bar{\alpha}_{t}:=\prod_{i=1}^{t}\alpha_{i}$. Eq. 1 is equivalent to optimizing a reweighted variational bound on negative log-likelihood. Without loss of generality, in this work, we parameterize $\bm{\epsilon}_{\theta}$ as the negative gradient of the energy function $\nabla_{\bm{x}_{t}}E_{\theta}(\bm{x}_{t},t)$ as in du2023reduce . To improve the energy landscape, du2024learning introduced a contrastive loss
@@
where ${\bm{x}}_{0}\sim p({\bm{x}})$, ${\bm{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, ${\bm{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$. ${\bm{x}}_{t}^{-}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}^{-}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$. Here ${\bm{x}}_{0}^{-}\sim p_{\text{corrupt}}({\bm{x}}_{0}^{-}|{\bm{x}}_{0})$ are negative examples by corrupting the positive examples ${\bm{x}}_{0}$. This contrastive loss encourages the positive (ground-truth) examples to have global energy minima.
The reverse process starts with $\bm{x}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$, and iteratively applies the learned denoising net $\bm{\epsilon}_{\theta}$, where song2020denoising introduces an adjustable noise scale $\sigma_{t}$:
@@
Importantly, song2020denoising highlights that a diffusion model trained with $\{\alpha_{t}\}_{t=1}^{T}$ can be used to sample with $\{\alpha_{\tau_{s}}\}_{s=1}^{S}$, where $\tau$ is any increasing sub-sequence of $1,2,\cdots,T$ of length $S$. This variable spacing significantly accelerates sampling and will serve as an efficient approach for MCTS simulation in Section 4.3.
@@ 4.1 Problem setup
In this work, we leverage an energy-based diffusion model to learn an energy function $E_{\theta}({\bm{c}},{\bm{x}}_{t},t)$ [^3] For simplicity, we omit ${\bm{c}}$ in subsequent notations. to model the optimization objective $\mathcal{J}({\bm{c}},{\bm{x}})$. During inference, we apply test-time scaling by varying the number of samples and use the learned energy function as an intrinsic verifier, thereby eliminating the need for an external verifier. However, the reasoning performance gain by test-time scaling is highly dependent on both the quality of the energy landscape and the inference-time search algorithm. To address these challenges, we propose VFScale as shown in Fig. 2, with our solutions detailed in Section 4.2 and Section 4.3.
@@ 4.2 Training of VFScale
Although the denoising loss ${\mathcal{L}}_{\text{MSE}}$ (Eq. 1) and the contrastive loss ${\mathcal{L}}_{\text{Contrast}}$ (Eq. 2) are effective, we find that simply scaling up inference budget through BoN processes yields marginal gain (Sec. 5.1). Through deeper analysis, we find that the core reason is the low quality of the energy landscape, espacially the lack of performance-energy consistency, which we address by introducing a novel Linear-Regression Negative Contrastive Learning loss and incorporating a KL regularization, which we detail as follows.
@@
Linear-Regression Negative Contrastive Learning. Specifically, while the contrastive loss drives positive examples to the global energy minimum, it imposes no constraints on the relative energy ordering among negative samples. For example, it is likely that a negative example ${\bm{x}}_{t}^{--}$ that is further apart from the positive example ${\bm{x}}_{t}$ has lower energy than a negative example ${\bm{x}}_{t}^{-}$ that is nearer. Thus, the inference process can be stuck around ${\bm{x}}_{t}^{--}$ and can hardly move out. This will result in reduced performance-energy consistency, where lower energy at step $t$ does not necessarily correspond to a more accurate predicted solution $\hat{{\bm{x}}}_{0}$, as will be shown in Sec. 5.1.
@@
This problem can not be easily remedied by increasing inference budget through BoN during test; instead, we explore a more fundamental way to solve this problem to shape the energy landscape by regularizing the energy among negative examples during training. The approach seeks to enforce consistency between the relative energy levels of negative samples and their corresponding performance quality, comprising the following key steps:
(1) Generate negative samples: Specifically, from a positive example ${\bm{x}}_{0}\sim p({\bm{x}})$, we sample two negative examples ${\bm{x}}_{0}^{-}$ and ${\bm{x}}_{0}^{--}$, [^4] Details to generate negative samples are in Appendix A. the latter has a larger L2 distance to ${\bm{x}}_{0}$. Specifically, the distance of ${\bm{x}}_{0}$, ${\bm{x}}_{0}^{-}$ and ${\bm{x}}_{0}^{--}$ to ${\bm{x}}_{0}$ are 00, $l_{2,0}^{-}=||{\bm{x}}_{0}^{-}-{\bm{x}}_{0}||_{2}^{2}$, and $l_{2,0}^{--}=||{\bm{x}}_{0}^{--}-{\bm{x}}_{0}||_{2}^{2}$, respectively.
@@
(2) Obtain the energy for each sample: Then we obtain their corresponding noisy examples at step $t$ via ${\bm{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$, ${\bm{x}}_{t}^{-}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}^{-}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$, and ${\bm{x}}_{t}^{--}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}^{--}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$. Their energy are $E^{+}_{t}=E_{\theta}({\bm{x}}_{t},t)$, $E^{-}_{t}=E_{\theta}({\bm{x}}_{t}^{-},t)$, and $E^{--}_{t}=E_{\theta}({\bm{x}}_{t}^{--},t)$, respectively.
(3) Apply linear regression to get the slope and intercept: To encourage a landscape where the energy difference between negative samples and the positive sample becomes smaller as they approach the positive sample, we directly constrain the relationship between the energy of different negative samples and the positive sample using linear regression. Specifically, we first apply the linear regression algorithm [^5] The details of linear regression algorithm can be found in Appendix A. to fit linear function through the three points $\{(0,E_{t}^{+}),(l_{2,0}^{-},E_{t}^{-}),(l_{2,0}^{--},E_{t}^{--})\}$, which is characterized by the slope $k_{t}$ and the intercept $b_{t}$.
@@
(4) Calculate the LRNCL loss: Then, we obtain the corresponding fitted points $\{(0,\hat{E}_{t}^{+}),(l_{2,0}^{-},\hat{E}_{t}^{-}),(l_{2,0}^{--},\hat{E}_{t}^{--})\}$ from the linear function. We then compute the Linear-Regression Negative Contrastive Learning (LRNCL) loss as:
${\mathcal{L}}_{\text{LRNCL}}={\mathbb{E}}_{{\bm{x}}_{0},{\bm{x}}_{0}^{-},{\bm{x}}_{0}^{--},{\bm{\epsilon}},t}\big{[}\max(0,\gamma-k_{t})\,+||E_{t}^{+}-\hat{E}_{t}^{+}||_{2}^{2}+||E_{t}^{-}-\hat{E}_{t}^{-}||_{2}^{2}+||E_{t}^{--}-\hat{E}_{t}^{--}||_{2}^{2}\big{]},$ (4)
@@
where $\gamma$ is a hyperparameter and $t\sim\{0,\dots,T\}$. The first term $\max(0,\gamma-k_{t})$ encourages that any three samples ${\bm{x}}_{t}$, ${\bm{x}}_{t}^{-}$, ${\bm{x}}_{t}^{--}$ have a positive (and larger-than-$\gamma$) slope of energy vs. L2 distance. The latter three terms encourage that the energy vs. L2 distance is linear, making the energy landscape smoother. [^6] It is also possible to use more complex shapes than linear regression to regularize the energy landscape, which we leave for future work. Here we find that simple linear regression works well.
@@
KL regularization. In addition to ${\mathcal{L}}_{\text{LRNCL}}$ that improves performance-energy consistency, we further include KL-regularization du2021improved to further improve the energy landscape:
@@
where $p_{\theta,t}({\bm{x}})$ is the probability distribution of ${\bm{x}}_{t}$ at denoising step $t$. When optimizing w.r.t. ${\mathcal{L}}_{\text{KL}}$, it is essentially optimizing w.r.t. the sampling (denoising) process by shaping the energy landscape to make it easier to sample. The first term in ${\mathcal{L}}_{\text{KL}}$ encourages the samples ${\bm{x}}_{t}$ to have low energy, and the second term is maximizing the entropy of the samples, encouraging the samples to be diverse. Both terms allow better test-time scaling. Different from du2021improved , we have this KL regularization on each denoising step $t$, and we employ a more accurate estimation of the entropy lombardi2016nonparametric .
Taken together, the training objective of our VFScale is:
${\mathcal{L}}={\mathcal{L}}_{\text{MSE}}+{\mathcal{L}}_{\text{Contrast}}+{\mathcal{L}}_{\text{LRNCL}}+{\mathcal{L}}_{\text{KL}},$ (6)

### v4

@@ Abstract
Inspired by human SYSTEM 2 thinking, LLMs excel at complex reasoning tasks via extended Chain-of-Thought. However, similar test-time scaling for diffusion models to tackle complex reasoning remains largely unexplored. From existing work, two primary challenges emerge in this setting: (i) the dependence on an external verifier indicating a notable gap from intrinsic reasoning of human intelligence without any external feedback, and (ii) the lack of an efficient search algorithm. In this paper, we introduce the Verifier-free Test-time Scalable Diffusion Model (VFScale) to achieve scalable intrinsic reasoning, which equips number-of-sample test-time scaling with the intrinsic energy function of diffusion models as the verifier. Concretely, VFScale comprises two key innovations to address the aforementioned challenges. On the training side, VFScale consists of a novel MRNCL loss and a KL regularization to improve the energy landscape, ensuring that the learned energy function itself serves as a reliable verifier. On the inference side, VFScale integrates the denoising process with a novel hybrid Monte Carlo Tree Search (hMCTS) to improve search efficiency. On challenging reasoning tasks of Maze and Sudoku, we demonstrate the effectiveness of VFScale’s training objective and scalable inference method. In particular, trained with Maze sizes of up to $6\times 6$, our VFScale solves 88% of Maze problems with much larger sizes of $15\times 15$, while standard diffusion models completely fail. The code can be found at https://github.com/AI4Science-WestlakeU/VFScale.
@@ 1 Introduction
Human intelligence can allocate more computation to solve harder problems, known as SYSTEM 2 thinking (Kahneman, 2011). Motivated by this, large language models (LLMs) based on autoregressive models have exhibited promising performance in reasoning tasks through deliberate long Chain-of-Thought (Lightman et al., 2023; DeepSeek-AI et al., 2025). Alongside autoregressive models, diffusion models have recently emerged as a compelling alternative. By casting reasoning as an optimization problem, these models iteratively refine candidate solutions toward higher-quality outputs. This approach has demonstrated strong potential, achieving notable results on tasks such as Sudoku solving, graph connectivity, and shortest-path computation (Du et al., 2022; 2024).
@@
Although the iterative denoising procedure of diffusion models mirrors the repeated reasoning cycles required by complex reasoning tasks, their performance declines sharply once the problem difficulty exceeds the training distribution (Du et al., 2022). Moreover, prior works show that simply increasing the number of sampling steps quickly leads to a performance plateau (Du et al., 2024). Instead, in image generation, inference‑time scaling to improve performance by increasing the number of samples has proven effective (Ma et al., 2025). This success, however, heavily relies on a dense external verifier (Ouyang et al., 2022; Lee et al., 2024), which provides continuous guidance by scoring sample quality. Moreover, these external verifiers are often learned from extensive labeled datasets that are particularly difficult to obtain for reasoning tasks. In contrast, human intelligence can engage in deep reflective intrinsic reasoning without external feedback, revealing a clear gap with such methods (Lombrozo, 2024; Chi et al., 1994; Huang et al., 2024). These observations motivate our central question: Can we design a Verifier‑free Test‑time Scalable Diffusion Model to achieve scalable intrinsic reasoning?
@@
In pursuit of such a test-time scalable intrinsic reasoning model, we leverage the fundamental mechanics of diffusion models. Diffusion models approximate the score function (Song et al., 2021b), which is equivalent to the negative gradient of the energy function. Our key insight is that the energy function, as a measure of the learned probability, can serve as an effective verifier itself. Based on the above premise, in this work, we explore equipping number-of-sample test-time scaling with diffusion models’ intrinsic energy function.
@@
However, this strategy presents two primary challenges: (1) Firstly, using number-of-sample test-time scaling requires a verifier to evaluate and select samples. Prior work on test-time scaling has also established that verifier accuracy is crucial for performance (Setlur et al., 2025; Ma et al., 2025). Consequently, using the energy function as an intrinsic verifier requires it to reliably reflect the sample quality. Nevertheless, previous energy-based diffusion models have largely ignored this requirement. (2) Secondly, the efficiency of the search algorithm is critical when scaling by varying the number of samples (Snell et al., 2025). However, the search efficiency of existing methods, such as best-of-$N$ (BoN), warrants further enhancement.
@@
To approach these challenges, we propose Verifier-Free Test-Time Scalable Diffusion Models (VFScale), a novel framework for solving complex reasoning problems. VFScale consists of innovations in both training objectives and inference scaling as shown in Fig. 2 to respectively address the two aforementioned challenges. (1) On the training side, VFScale incorporates a novel Monotonic-Regression Negative Contrastive Learning (MRNCL) objective and a KL regularization to improve the energy landscape, especially to better align the energy function with sample quality with MRNCL loss. The MRNCL objective samples two negatives [^1] Here, positive samples refer to data points, while negative samples correspond to noise-perturbed versions of the data. at different corruption levels for each positive sample, and requires their energy values to reflect their L2 distances to the positive. We refer to this alignment as performance-energy consistency. [^2] Detailed description for performance-energy consistency can be found in Appendix A. The KL regularization term backpropagates through the denoising process, smoothing the energy landscape. (2) On the inference side, we propose a novel hybrid Monte Carlo Tree Search (hMCTS) that strategically balances exploration and exploitation during the denoising process. In the noisy initial stages, it employs Best-of-N (BoN) for broad, parallel exploration, preventing the premature elimination of promising solution pathways. As denoising progresses and the search space narrows, it transitions to MCTS for deep, fine-grained exploitation of the most viable candidates. This hybrid strategy enables VFScale to efficiently translate increased test-time computation into substantial gains in reasoning performance.
@@
In summary, our contributions are as follows: (1) We introduce Verifier-free Test-time Scalable Diffusion Models (VFScale), a novel framework that can scale up test-time computation for better reasoning capability. (2) We introduce Monotonic-Regression Negative Contrastive Learning (MRNCL) and KL regularization into the training objective to improve the energy landscape. (3) We integrate a hybrid Monte Carlo Tree Search into the denoising process, enabling test-time scaling with better search efficiency. (4) We demonstrate the effectiveness of our method on challenging Sudoku and Maze datasets where the conditions are out-of-distribution or the problem sizes are much larger.
Figure 2: Overview of VFScale. This figure illustrates the key aspects of VFScale by contrasting its training and inference strategies with those of the previous method. (1) To qualify the intrinsic energy of diffusion models as a verifier, VFScale introduces ${\mathcal{L}}_{\text{MRNCL}}$ and ${\mathcal{L}}_{\text{KL}}$ to improve the energy landscape during training. (2) In order for a higher search efficiency, VFScale proposes hybrid Monte Carlo Tree Search (hMCTS) that achieves a balance between best-of-NN and MCTS.
@@ 2 Related Work
Diffusion for Reasoning: The iterative refinement process inherent in diffusion models lends itself well to tasks requiring multistep reasoning. Du et al. (2022) firstly formulated multi-step reasoning as an energy minimization problem. To approach harder problems with more complex energy landscapes, more compute budget is used for optimization steps. Du et al. (2024) formulated the energy-based model as a diffusion model, and the performance on harder tasks was significantly improved. While effective, the performance gain by scaling the compute of these methods quickly saturates as the number of sampling steps increases. Building upon Du et al. (2024), VFScale instead explores verifier-free number-of-sample scaling and proposes corresponding training and inference strategies.
Test-time Scaling of Diffusion: Improving diffusion model samples by increasing compute resources at inference time has emerged as an increasingly important direction. Early approaches focused on increasing the number of sampling steps, leading to diminishing returns (Ho et al., 2020; Karras et al., 2022). More recent work (Ma et al., 2025; Yoon et al., 2025) has explored alternative scaling dimensions, such as increasing the number of sampled candidates, though these methods typically rely on external verifiers to guide solution selection. VFScale advances beyond these methods by leveraging the intrinsic energy function to guide test-time search, a capability further cultivated through customized training and inference strategies. These innovations make VFScale a versatile and scalable solution for real-world reasoning tasks.
@@ 3 Preliminary
The Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) contains a predefined forward process to corrupt data into Gaussian noises, and a learnable reverse process to generate new samples from them. The forward process follows a Gaussian transition kernel $q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t}|\sqrt{\alpha_{t}}\bm{x}_{t-1},(1-\alpha_{t})\bm{I}),t=1,\cdots,T$, where the noise schedule $\{\alpha_{t}\}_{t=1}^{T}$ and $T$ is set to make $\bm{x}_{T}$ follow approximately a standard Gaussian distribution. The reverse process can be learned to predict the noise from the corrupted data by minimizing
@@
where the expectation is w.r.t. $\bm{x}_{0}\sim p(\bm{x})$, $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$, and $t\sim\{1,...T\}$. $\bar{\alpha}_{t}:=\prod_{i=1}^{t}\alpha_{i}$. Eq. 1 is equivalent to optimizing a reweighted variational bound on negative log-likelihood. Without loss of generality, in this work, we parameterize $\bm{\epsilon}_{\theta}$ as the negative gradient of the energy function $\nabla_{\bm{x}_{t}}E_{\theta}(\bm{x}_{t},t)$ as in Du et al. (2023). To improve the energy landscape, Du et al. (2024) introduced a contrastive loss
@@
where ${\bm{x}}_{0}\sim p({\bm{x}})$, ${\bm{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, ${\bm{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$. ${\bm{x}}_{t}^{-}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}^{-}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$. Here ${\bm{x}}_{0}^{-}\sim p_{\text{corrupt}}({\bm{x}}_{0}^{-}|{\bm{x}}_{0})$ are negative examples by corrupting the positive examples ${\bm{x}}_{0}$. This contrastive loss encourages the positive (ground-truth) examples to be local energy minima.
The reverse process starts with $\bm{x}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$, and iteratively applies the learned denoising net $\bm{\epsilon}_{\theta}$, where Song et al. (2021a) introduces an adjustable noise scale $\sigma_{t}$:
@@
Importantly, Song et al. (2021a) highlights that a diffusion model trained with $\{\alpha_{t}\}_{t=1}^{T}$ can be used to sample with $\{\alpha_{\tau_{s}}\}_{s=1}^{S}$, where $\tau$ is any increasing subsequence of $1,2,\cdots,T$ of length $S$. This variable spacing significantly accelerates sampling and will serve as an efficient approach for MCTS simulation in Section 4.3.
@@ 4.1 Problem setup
In this work, we leverage an energy-based diffusion model to learn an energy function $E_{\theta}({\bm{c}},{\bm{x}}_{t},t)$ to model the optimization objective $\mathcal{J}({\bm{c}},{\bm{x}})$. For the sake of brevity, we omit $\bm{c}$} in subsequent notations and denote it simply as $E_{\theta}({\bm{x}}_{t},t)$. During inference, we apply test-time scaling by varying the number of samples and use the learned energy function as an intrinsic verifier, thereby eliminating the need for an external verifier. However, the reasoning performance gain by test-time scaling is highly dependent on both the quality of the energy landscape and the inference-time search algorithm. To address these challenges, we propose VFScale as shown in Fig. 2, with our solutions detailed in Section 4.2 and Section 4.3.
@@ 4.2 Training of VFScale
Although the denoising loss ${\mathcal{L}}_{\text{MSE}}$ (Eq. 1) and the contrastive loss ${\mathcal{L}}_{\text{Contrast}}$ (Eq. 2) are effective, we find that simply scaling up inference budget through BoN processes yields only a marginal gain (Sec. 5.1). Through a deeper analysis, we find that the core reason is the low quality of the energy landscape, especially the lack of performance-energy consistency, which we address by introducing a novel Monotonic-Regression Negative Contrastive Learning (MRNCL) loss and incorporating a KL regularization, which we detail as follows.
@@
Monotonic-Regression Negative Contrastive Learning. Specifically, while the contrastive loss drives positive examples to the local energy minimum, it imposes no constraints on the relative energy ordering among negative samples. For example, it is likely that a negative example ${\bm{x}}_{t}^{--}$ that is further apart from the positive example ${\bm{x}}_{t}$ has lower energy (higher probability) than a negative example ${\bm{x}}_{t}^{-}$ that is nearer. Thus, the inference process can be stuck around ${\bm{x}}_{t}^{--}$ and can hardly move out. This will result in reduced performance-energy consistency, where a lower energy at step $t$ does not necessarily correspond to a more accurate predicted solution $\hat{{\bm{x}}}_{0}$, as will be shown in Sec. 5.1.
@@
This problem cannot be easily remedied by increasing inference budget through BoN during test; instead, we explore a more fundamental way to solve this problem to shape the energy landscape by regularizing the energy among negative examples during training. The approach seeks to enforce consistency between the relative energy levels of negative samples and their corresponding performance quality, comprising the following key steps.
[…]
@@
(2) Obtain the energy for each sample: Then we obtain their corresponding noisy examples at step $t$ via ${\bm{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$, ${\bm{x}}_{t}^{-}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}^{-}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$, and ${\bm{x}}_{t}^{--}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}^{--}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$. Their energy values are $E^{+}_{t}=E_{\theta}({\bm{x}}_{t},t)$, $E^{-}_{t}=E_{\theta}({\bm{x}}_{t}^{-},t)$, and $E^{--}_{t}=E_{\theta}({\bm{x}}_{t}^{--},t)$, respectively.
(3) Apply Monotonic regression to get the slope and intercept: The core idea is to directly constrain the relationship between the energy of different samples and their distance to the ground truth using a monotonic function. Linear regression is one specific and effective instance of this approach [^4] For its simplicity and effectiveness, we primarily use linear regression in this work and term the specific loss LRNCL (Linear-Regression NCL). We provide a detailed ablation study in Appendix D.2 that analyzes the impact of other monotonic constraints (e.g., quadratic, rank-based) and further justifies the choice of a linear function for achieving optimal test-time scalability., which will be used throughout the paper unless otherwise mentioned. Specifically, we fit a linear function [^5] The details of linear regression algorithm can be found in Appendix A. with the three points $\{(0,E_{t}^{+}),(l_{2,0}^{-},E_{t}^{-}),(l_{2,0}^{--},E_{t}^{--})\}$, which is characterized by the slope $k_{t}$ and the intercept $b_{t}$.
@@
(4) Calculate the MRNCL loss: Then we obtain the corresponding fitted points $\{(0,\hat{E}_{t}^{+}),(l_{2,0}^{-},\hat{E}_{t}^{-}),(l_{2,0}^{--},\hat{E}_{t}^{--})\}$ from the linear function. We then compute the MRNCL loss as follows:
[…]
@@
where $\gamma$ is a hyperparameter and $t\sim\{0,\dots,T\}$. The first term $\max(0,\gamma-k_{t})$ encourages that any three samples ${\bm{x}}_{t}$, ${\bm{x}}_{t}^{-}$, ${\bm{x}}_{t}^{--}$ have a positive (and larger than $\gamma$) slope of energy vs. L2 distance. The latter three terms encourage the energy vs. L2 distance to be linear, making the energy landscape smoother.
@@
KL regularization. In addition to ${\mathcal{L}}_{\text{MRNCL}}$ that improves performance-energy consistency, we further include KL-regularization (Du et al., 2021) to further improve the energy landscape:
@@
where $p_{\theta,t}({\bm{x}})$ is the probability distribution of ${\bm{x}}_{t}$ at denoising step $t$. When optimizing w.r.t. ${\mathcal{L}}_{\text{KL}}$, it is essentially optimizing w.r.t. the sampling (denoising) process by shaping the energy landscape to make it easier to sample. The first term in ${\mathcal{L}}_{\text{KL}}$ encourages the samples ${\bm{x}}_{t}$ to have low energy, and the second term is maximizing the entropy of the samples, encouraging the samples to be diverse. Both terms allow better test-time scaling. Unlike Du et al. (2021), we have this KL regularization on each denoising step $t$, and we employ a more accurate estimation of entropy (Lombardi & Pant, 2016).
Together, the training objective of our VFScale is:
${\mathcal{L}}={\mathcal{L}}_{\text{MSE}}+{\mathcal{L}}_{\text{Contrast}}+{\mathcal{L}}_{\text{MRNCL}}+{\mathcal{L}}_{\text{KL}},$ (6)